OpenStudy (gracefulchibi):

PLEASE HELP!! The average value of a certain type of automobile was $13,320 in 1995 and depreciated to $4020 in 2000. Let y be the average value of the automobile in the year x, where x = 0 represents 1995. Write a linear equation that models the value of the automobile in terms of the year x.

OpenStudy (gracefulchibi):

@jigglypuff314 please help me~ T^T

OpenStudy (helpkay):

I will try to help but Im not 100% sure of what you are doing

OpenStudy (gracefulchibi):

thank you! the answer options are : y=-1860x-13320 y=-1860x+4020 y=-1/1860x-4020 y=-1860x-5280

OpenStudy (gracefulchibi):

@helpkay I think its slope/intercept y=mx+b

OpenStudy (helpkay):

yes that is how you find the slope but i do not get the -1860

OpenStudy (gracefulchibi):

right~ I don't know where that or the 5280 came from

OpenStudy (helpkay):

I will look it up but I don't know how to solve this problem

OpenStudy (gracefulchibi):

ok thanks c:

OpenStudy (helpkay):

http://openstudy.com/updates/5077c8e8e4b02f109be3ff60 now that I see how to do it it was not that hard at all